A man is two times older than his son. together their ages equal 63, what are their ages?
Answer by Alan3354(69443)

A man is two times older than his son. together their ages equal 63, what are their ages?
====================
two times older = 3 times as old
---
S = son's age
M = man's age
---
M = 3S
M + S = 63
3S + S = 63
4S = 63
Son's age = 63/4 = 15.75
Man's age = 47.25
=============================
2 times older is 200% older.
Age + 200% of Age = 300%
= 3 times as old.
-----------------
2 times older is not the same as 2 times as old.
A lot of people don't understand that, and some of those write math books, unfortunately.